# Training Large Language Models to Reason in a Continuous Latent Space

### v2

@@ Abstract
Large language models (LLMs) are restricted to reason in the “language space”, where they typically express the reasoning process with a chain-of-thought (CoT) to solve a complex reasoning problem. However, we argue that language space may not always be optimal for reasoning. For example, most word tokens are primarily for textual coherence and not essential for reasoning, while some critical tokens require complex planning and pose huge challenges to LLMs. To explore the potential of LLM reasoning in an unrestricted latent space instead of using natural language, we introduce a new paradigm Coconut (Chain of Continuous Thought). We utilize the last hidden state of the LLM as a representation of the reasoning state (termed “continuous thought”). Rather than decoding this into a word token, we feed it back to the LLM as the subsequent input embedding directly in the continuous space. Experiments show that Coconut can effectively augment the LLM on several reasoning tasks. This novel latent reasoning paradigm leads to emergent advanced reasoning patterns: the continuous thought can encode multiple alternative next reasoning steps, allowing the model to perform a breadth-first search (BFS) to solve the problem, rather than prematurely committing to a single deterministic path like CoT. Coconut outperforms CoT in certain logical reasoning tasks that require substantial backtracking during planning, with fewer thinking tokens during inference. These findings demonstrate the promise of latent reasoning and offer valuable insights for future research.
@@ 1 Introduction
A significant issue arises when LLMs use language for reasoning: the amount of reasoning required for each particular reasoning token varies greatly, yet current LLM architectures allocate nearly the same computing budget for predicting every token. Most tokens in a reasoning chain are generated solely for fluency, contributing little to the actual reasoning process. On the contrary, some critical tokens require complex planning and pose huge challenges to LLMs. While previous work has attempted to fix these problems by prompting LLMs to generate succinct reasoning chains (Madaan and Yazdanbakhsh, 2022), or performing additional reasoning before generating some critical tokens (Zelikman et al., 2024), these solutions remain constrained within the language space and do not solve the fundamental problems. On the contrary, it would be ideal for LLMs to have the freedom to reason without any language constraints, and then translate their findings into language only when necessary.
@@
Experimentally, Coconut successfully enhances the reasoning capabilities of LLMs. For math reasoning (GSM8k, Cobbe et al., 2021), using continuous thoughts is shown to be beneficial to reasoning accuracy, mirroring the effects of language reasoning chains. This indicates the potential to scale and solve increasingly challenging problems by chaining more continuous thoughts. On logical reasoning including ProntoQA (Saparov and He, 2022), and our newly proposed ProsQA (Section 4.1) which requires stronger planning ability, Coconut and some of its variants even surpasses language-based CoT methods, while generating significantly fewer tokens during inference. We believe that these findings underscore the potential of latent reasoning and could provide valuable insights for future research.
@@ 2 Related Work
Chain-of-thought (CoT) reasoning. We use the term chain-of-thought broadly to refer to methods that generate an intermediate reasoning process in language before outputting the final answer. This includes prompting LLMs (Wei et al., 2022; Khot et al., 2022; Zhou et al., 2022), or training LLMs to generate reasoning chains, either with supervised finetuning (Yue et al., 2023; Yu et al., 2023) or reinforcement learning (Wang et al., 2024; Havrilla et al., 2024; Shao et al., 2024; Yu et al., 2024a). Madaan and Yazdanbakhsh (2022) classified the tokens in CoT into symbols, patterns, and text, and proposed to guide the LLM to generate concise CoT based on analysis of their roles. Recent theoretical analyses have demonstrated the usefulness of CoT from the perspective of model expressivity (Feng et al., 2023; Merrill and Sabharwal, 2023; Li et al., 2024). By employing CoT, the effective depth of the transformer increases because the generated outputs are looped back to the input (Feng et al., 2023). These analyses, combined with the established effectiveness of CoT, motivated our design that feeds the continuous thoughts back to the LLM as the next input embedding. While CoT has proven effective for certain tasks, its autoregressive generation nature makes it challenging to mimic human reasoning on more complex problems (LeCun, 2022; Hao et al., 2023), which typically require planning and search. There are works that equip LLMs with explicit tree search algorithms (Xie et al., 2023; Yao et al., 2023; Hao et al., 2024), or train the LLM on search dynamics and trajectories (Lehnert et al., 2024; Gandhi et al., 2024; Su et al., 2024). In our analysis, we find that after removing the constraint of a language space, a new reasoning pattern similar to BFS emerges, even though the model is not explicitly trained in this way.
@@
Latent reasoning in LLMs. Previous works mostly define latent reasoning in LLMs as the hidden computation in transformers (Yang et al., 2024; Biran et al., 2024). Yang et al. (2024) constructed a dataset of two-hop reasoning problems and discovered that it is possible to recover the intermediate variable from the hidden representations. Biran et al. (2024) further proposed to intervene the latent reasoning by “back-patching” the hidden representation. Shalev et al. (2024) discovered parallel latent reasoning paths in LLMs. Another line of work has discovered that, even if the model generates a CoT to reason, the model may actually utilize a different latent reasoning process. This phenomenon is known as the unfaithfulness of CoT reasoning (Wang et al., 2022; Turpin et al., 2024). To enhance the latent reasoning of LLM, previous research proposed to augment it with additional tokens. Goyal et al. (2023) pretrained the model by randomly inserting a learnable <pause> tokens to the training corpus. This improves LLM’s performance on a variety of tasks, especially when followed by supervised finetuning with <pause> tokens. On the other hand, Pfau et al. (2024) further explored the usage of filler tokens, e.g., “...”, and concluded that they work well for highly parallelizable problems. However, Pfau et al. (2024) mentioned these methods do not extend the expressivity of the LLM like CoT; hence, they may not scale to more general and complex reasoning problems. Wang et al. (2023) proposed to predict a planning token as a discrete latent variable before generating the next reasoning step. Recently, it has also been found that one can “internalize” the CoT reasoning into latent reasoning in the transformer with knowledge distillation (Deng et al., 2023) or a special training curriculum which gradually shortens CoT (Deng et al., 2024). Yu et al. (2024b) also proposed to distill a model that can reason latently from data generated with complex reasoning algorithms. These training methods can be combined to our framework, and specifically, we find that breaking down the learning of continuous thoughts into multiple stages, inspired by iCoT (Deng et al., 2024), is very beneficial for the training. Recently, looped transformers (Giannou et al., 2023; Fan et al., 2024) have been proposed to solve algorithmic tasks, which have some similarities to the computing process of continuous thoughts, but we focus on common reasoning tasks and aim at investigating latent reasoning in comparison to language space.
@@ 3 Coconut: Chain of Continuous Thought
4 Experiments
We validate the feasibility of LLM reasoning in a continuous latent space through experiments on three datasets. We mainly evaluate the accuracy by comparing the model-generated answers with the ground truth. The number of newly generated tokens per question is also analyzed, as a measure of reasoning efficiency. We report the clock-time comparison in Appendix 8.
4.1 Reasoning Tasks
Math Reasoning. We use GSM8k (Cobbe et al., 2021) as the dataset for math reasoning. It consists of grade school-level math problems. Compared to the other datasets in our experiments, the problems are more diverse and open-domain, closely resembling real-world use cases. Through this task, we explore the potential of latent reasoning in practical applications. To train the model, we use a synthetic dataset generated by Deng et al. (2023).
Logical Reasoning. Logical reasoning involves the proper application of known conditions to prove or disprove a conclusion using logical rules. This requires the model to choose from multiple possible reasoning paths, where the correct decision often relies on exploration and planning ahead. We use 5-hop ProntoQA (Saparov and He, 2022) questions, with fictional concept names. For each problem, a tree-structured ontology is randomly generated and described in natural language as a set of known conditions. The model is asked to judge whether a given statement is correct based on these conditions. This serves as a simplified simulation of more advanced reasoning tasks, such as automated theorem proving (Chen et al., 2023; DeepMind, 2024).
We found that the generation process of ProntoQA could be more challenging, especially since the size of distracting branches in the ontology is always small, reducing the need for complex planning. To fix that, we apply a new dataset construction pipeline using randomly generated DAGs to structure the known conditions. The resulting dataset requires the model to perform substantial planning and searching over the graph to find the correct reasoning chain. We refer to this new dataset as ProsQA (Proof with Search Question-Answering). A visualized example is shown in Figure 7. More details of datasets can be found in Appendix 7.
@@ 4.2 Experimental Setup
We use a pre-trained GPT-2 (Radford et al., 2019) as the base model for all experiments. The learning rate is set to $1\times 10^{-4}$ while the effective batch size is 128. Following Deng et al. (2024), we also reset the optimizer when the training stages switch.
Math Reasoning. By default, we use 2 latent thoughts (i.e., $c=2$) for each reasoning step. We analyze the correlation between performance and $c$ in Section 1. The model goes through 3 stages besides the initial stage. Then, we have an additional stage, where we still use $3\times c$ continuous thoughts as in the penultimate stage, but remove all the remaining language reasoning chain. This handles the long-tail distribution of reasoning chains longer than 3 steps. We train the model for 6 epochs in the initial stage, and 3 epochs in each remaining stage.
Logical Reasoning. We use one continuous thought for every reasoning step (i.e., $c=1$). The model goes through 6 training stages in addition to the initial stage, because the maximum number of reasoning steps is 6 in these two datasets. The model then fully reasons with continuous thoughts to solve the problems in the last stage. We train the model for 5 epochs per stage.
For all datasets, after the standard schedule, the model stays in the final training stage, until the 50th epoch. We select the checkpoint based on the accuracy on the validation set. For inference, we manually set the number of continuous thoughts to be consistent with their final training stage. We use greedy decoding for all experiments.
4.3 Baselines and Variants of Coconut
We consider the following baselines: (1) CoT: We use the complete reasoning chains to train the language model with supervised finetuning, and during inference, the model generates a reasoning chain before outputting an answer. (2) No-CoT: The LLM is trained to directly generate the answer without using a reasoning chain. (3) iCoT (Deng et al., 2024): The model is trained with language reasoning chains and follows a carefully designed schedule that “internalizes” CoT. As the training goes on, tokens at the beginning of the reasoning chain are gradually removed until only the answer remains. During inference, the model directly predicts the answer. (4) Pause token (Goyal et al., 2023): The model is trained using only the question and answer, without a reasoning chain. However, different from No-CoT, special <pause> tokens are inserted between the question and answer, which are believed to provide the model with additional computational capacity to derive the answer. For a fair comparison, the number of <pause> tokens is set the same as continuous thoughts in Coconut.
We also evaluate some variants of our method: (1) w/o curriculum: Instead of the multi-stage training, we directly use the data from the last stage which only includes questions and answers to train Coconut. The model uses continuous thoughts to solve the whole problem. (2) w/o thought: We keep the multi-stage training which removes language reasoning steps gradually, but don’t use any continuous latent thoughts. While this is similar to iCoT in the high-level idea, the exact training schedule is set to be consistent with Coconut, instead of iCoT. This ensures a more strict comparison. (3) Pause as thought: We use special <pause> tokens to replace the continuous thoughts, and apply the same multi-stage training curriculum as Coconut.
4.4 Results and Discussion

### v3

@@ Abstract
Large language models (LLMs) are restricted to reason in the “language space”, where they typically express the reasoning process with a chain-of-thought (CoT) to solve a complex reasoning problem. However, we argue that language space may not always be optimal for reasoning. For example, most word tokens primarily ensure textual coherence and are not essential for reasoning, while some critical tokens require complex planning and pose huge challenges to LLMs. To explore the potential of LLM reasoning in an unrestricted latent space instead of using natural language, we introduce a new paradigm Coconut (Chain of Continuous Thought). We utilize the last hidden state of the LLM as a representation of the reasoning state (termed “continuous thought”). Rather than decoding this into a word token, we feed it back to the LLM as the subsequent input embedding directly in the continuous space. This latent reasoning paradigm leads to the emergence of an advanced reasoning pattern: the continuous thought can encode multiple alternative next reasoning steps, allowing the model to perform a breadth-first search (BFS) to solve the problem, rather than prematurely committing to a single deterministic path like CoT. Coconut outperforms CoT on certain logical reasoning tasks that require substantial search during planning, and shows a better trade-off between accuracy and efficiency.
@@ 1 Introduction
A significant issue arises when LLMs use language for reasoning: the amount of reasoning required for each particular token varies greatly, yet current LLM architectures allocate nearly the same computing budget for predicting every token. Most tokens in a reasoning chain are generated solely for fluency, contributing little to the actual reasoning process. By contrast, some critical tokens require complex planning and pose huge challenges to LLMs. While previous work has attempted to fix these problems by prompting LLMs to generate succinct reasoning chains (Madaan and Yazdanbakhsh, 2022), or performing additional reasoning before generating some critical tokens (Zelikman et al., 2024), these solutions remain constrained within the language space and do not solve the fundamental problems. On the contrary, it would be ideal for LLMs to have the freedom to reason without any language constraints, and then translate their findings into language only when necessary.
@@
Experimentally, Coconut successfully enhances the reasoning capabilities of LLMs. For math reasoning (GSM8k, Cobbe et al., 2021), using continuous thoughts is shown to be beneficial to reasoning accuracy, mirroring the effects of language reasoning chains. This indicates the potential to scale and solve increasingly challenging problems by chaining more continuous thoughts. On logical reasoning including ProntoQA (Saparov and He, 2022), and our newly proposed ProsQA (Section 4) which requires stronger planning ability, Coconut and some of its variants even surpasses language-based CoT methods, while generating significantly fewer tokens during inference. We believe that these findings underscore the potential of latent reasoning and could provide valuable insights for future research.
@@ 2 Related Work
Chain-of-thought (CoT) reasoning. We use the term chain-of-thought broadly to refer to methods that generate an intermediate reasoning process in language before outputting the final answer. This includes prompting LLMs (Wei et al., 2022; Khot et al., 2022; Zhou et al., 2022), or training LLMs to generate reasoning chains, either with supervised finetuning (Yue et al., 2023; Yu et al., 2023) or reinforcement learning (Wang et al., 2024; Havrilla et al., 2024; Shao et al., 2024; Yu et al., 2024a). Madaan and Yazdanbakhsh (2022) classified the tokens in CoT into symbols, patterns, and text, and proposed to guide the LLM to generate concise CoT based on analysis of their roles. Recent theoretical analyses have demonstrated the usefulness of CoT from the perspective of model expressivity (Feng et al., 2023; Merrill and Sabharwal, 2023; Li et al., 2024). By employing CoT, the effective depth of the transformer increases because the generated outputs are looped back to the input (Feng et al., 2023). These analyses, combined with the established effectiveness of CoT, motivated our design of feeding the continuous thoughts back into the LLM as input embeddings. While CoT has proven effective for certain tasks, its autoregressive generation nature makes it challenging to mimic human reasoning on more complex problems (LeCun, 2022; Hao et al., 2023), which typically require planning and search. There are works that equip LLMs with explicit tree search algorithms (Xie et al., 2023; Yao et al., 2023; Hao et al., 2024), or train the LLM on search dynamics and trajectories (Lehnert et al., 2024; Gandhi et al., 2024; Su et al., 2024). In our analysis, we find that after removing the constraint of a language space, a new reasoning pattern similar to BFS emerges, even though the model is not explicitly trained in this way.
@@
Latent reasoning in LLMs. Previous works mostly define latent reasoning in LLMs as the hidden computation in transformers (Yang et al., 2024; Biran et al., 2024). Yang et al. (2024) constructed a dataset of two-hop reasoning problems and discovered that it is possible to recover the intermediate variable from the hidden representations. Biran et al. (2024) further proposed to intervene the latent reasoning by “back-patching” the hidden representation. Shalev et al. (2024) discovered parallel latent reasoning paths in LLMs. Another line of work has discovered that, even if the model generates a CoT to reason, the model may actually utilize a different latent reasoning process. This phenomenon is known as the unfaithfulness of CoT reasoning (Wang et al., 2022; Turpin et al., 2024). To enhance the latent reasoning of LLMs, previous research proposed to augment it with additional tokens. Goyal et al. (2023) pretrained the model by randomly inserting a learnable <pause> tokens to the training corpus. This improves LLM’s performance on a variety of tasks, especially when followed by supervised finetuning with <pause> tokens. On the other hand, Pfau et al. (2024) further explored the usage of filler tokens, e.g., “...”, and concluded that they work well for highly parallelizable problems. However, Pfau et al. (2024) mentioned these methods do not extend the expressivity of the LLM like CoT; hence, they may not scale to more general and complex reasoning problems. Wang et al. (2023) proposed to predict a planning token as a discrete latent variable before generating the next reasoning step. Recently, it has also been found that one can “internalize” the CoT reasoning into latent reasoning in the transformer with knowledge distillation (Deng et al., 2023) or a special training curriculum which gradually shortens CoT (Deng et al., 2024). Yu et al. (2024b) also proposed to distill a model that can reason latently from data generated with complex reasoning algorithms. These training methods can be combined to our framework, and specifically, we find that breaking down the learning of continuous thoughts into multiple stages, inspired by iCoT (Deng et al., 2024), is very beneficial for the training. Other work explores alternative architectures for latent reasoning, including looped transformers (Giannou et al., 2023; Fan et al., 2024), diffusion models in sentence embedding space (Barrault et al., 2024). Different from these works, we focus on general multi-step reasoning tasks and aim to investigate the unique properties of latent reasoning in comparison to language space. In addition to reasoning tasks, Pham et al. (2023) also explored using continuous space for multi-agent communication. Building on Coconut, Zhu et al. (2025b) developed a theoretical framework demonstrating that continuous CoT can be more efficient than discrete CoT on certain tasks by encoding multiple reasoning paths in superposition states. Subsequently, Zhu et al. (2025a) analyzed the training dynamics to explain how such superposition emerges under the Coconut training objective.
@@ 3 Coconut: Chain of Continuous Thought
4 Continuous Space Enables Latent Tree Search
In this section, we provide a proof of concept of the advantage of continuous latent space reasoning. On ProsQA, a new dataset that requires extensive planning ability, Coconut outperforms language space CoT reasoning. Interestingly, our analysis indicates that the continuous representation of reasoning can encode multiple alternative next reasoning steps. This allows the model to perform a breadth-first search (BFS) to solve the problem, instead of prematurely committing to a single deterministic path like language CoT.
We start by introducing the experimental setup (Section 4.1). By leveraging Coconut’s ability to switch between language and latent space reasoning, we are able to control the model to interpolate between fully latent reasoning and fully language reasoning and test their performance (Section 4.2). This also enables us to interpret the latent reasoning process as tree search (Section 4.3). Based on this perspective, we explain why latent reasoning can help LLMs make better decisions (Section 4.4).
@@ 4.1 Experimental Setup
Dataset. We introduce ProsQA (Proof with Search Question-Answering), a new logical reasoning dataset. A visualized example is shown in Figure 5. Each instance in ProsQA consists of a directed acyclic graph (DAG) of logical relationships between concepts, presented as natural language statements. The task requires models to determine logical relationships by finding valid paths through this graph, demanding sophisticated planning and search strategies. Unlike previous logical reasoning datasets like ProntoQA (Saparov and He, 2022), ProsQA’s DAG structure introduces complex exploration paths, making it particularly challenging for models to identify the correct reasoning chain. More comprehensive details about the dataset construction and characteristics can be found in Appendix 7.
Setup. We use a pre-trained GPT-2 model as the base model for all experiments. The learning rate is set to $1\times 10^{-4}$ while the effective batch size is 128. We train a Coconut model following the training procedure in Section 3. Since the maximum reasoning steps in ProsQA is 6, we set the number of training stages to $N=6$ in the training procedure. In each stage, we train the model for 5 epochs, and stay in the last stage until the 50 epochs. The checkpoint with the best accuracy in the last stage is used for evaluation. As reference, we report the performance of (1) CoT: the model is trained with CoT data, and during inference, the model will generate a complete reasoning chain to solve the problem. (2) no-CoT: the model is trained with only the question and answer pairs, without any reasoning steps. During inference, the model will output the final answer directly.
To understand the properties of latent and language reasoning space, we manipulate the model to switch between fully latent reasoning and fully language reasoning, by manually setting the position of the <eot> token during inference. When we enforce Coconut to use $k$ continuous thoughts, the model is expected to output the remaining reasoning chain in language, starting from the $k+1$ step. In our experiments, we test variants of Coconut on ProsQA with $k\in\{0,1,2,3,4,5,6\}$. Note that all these variants only differ in inference time while sharing the same model weights.
Metrics. We apply two sets of evaluation metrics. One of them is based on the correctness of the final answer, regardless of the reasoning process. It is also the main metric used in the later sections (Section 5.3). To enable fine-grained analysis on ProsQA, we define another metric on the reasoning process. We classify a reasoning chain into (1) Correct Path: The output is one of the shortest paths to the correct answer. (2) Longer Path: A valid path that correctly answers the question but is longer than the shortest path. (3) Hallucination: The path includes nonexistent edges or is disconnected. (4) Wrong Target: A valid path in the graph, but the destination node is not the one being asked. These four categories naturally apply to the output from Coconut ($k=0$) and CoT, which generate the full path. For Coconut with $k>0$ that outputs only partial paths in language (with the initial steps in continuous reasoning), we classify the reasoning as a Correct Path if a valid explanation can complete it. Also, we define Longer Path and Wrong Target for partial paths similarly. If no valid explanation completes the path, it’s classified as Hallucination. In no-CoT and Coconut with larger $k$, the model may only output the final answer without any partial path, and it falls into (5) Correct Label or (6) Incorrect Label. These six categories cover all cases without overlap.
4.2 Overall Results
Figure 3 presents a comparative analysis of various reasoning methods evaluated on ProsQA. The model trained using CoT frequently hallucinates non-existent edges or outputs paths leading to incorrect targets, resulting in lower answer accuracy. In contrast, Coconut, which leverages continuous space reasoning, demonstrates improved accuracy as it utilizes an increasing number of continuous thoughts. Additionally, the rate of correct reasoning processes (indicated by “Correct Label” and “Correct Path”) significantly increases. At the same time, there is a notable reduction in instances of “Hallucination” and “Wrong Target,” issues that typically emerge when the model makes mistakes early in the reasoning process.
An intuitive demonstration of the limitations of reasoning in language space is provided by the case study depicted in Figure 5. As shown, models operating in language space often fail to plan ahead or backtrack. Once they commit to an incorrect path, they either hallucinate unsupported edges or terminate with irrelevant conclusions. In contrast, latent reasoning avoids such premature commitments by enabling the model to iteratively refine its decisions across multiple reasoning steps. This flexibility allows the model to progressively eliminate incorrect options and converge on the correct answer, ultimately resulting in higher accuracy.
4.3 Interpreting the Latent Reasoning as Tree Search
To better understand Coconut, we probe the latent reasoning process by forcing the model to explicitly generate language reasoning steps following intermediate continuous thoughts (Figure 5). Using the example presented in Figure 5, at the initial reasoning step, the model must select which immediate child node of “Alex” to consider next, specifically from the set {“lempus”, “sterpus”, “zhorpus”, “grimpus”}. The distribution over these candidate next steps is visualized in Figure 5, left. In the subsequent reasoning step, these nodes expand further into an extended set of potential paths, including all grandchildren of “Alex” (Figure 5, right).
We define the predicted probability of a concept following continuous thoughts as a value function (Figure 5), estimating each node’s potential for reaching the correct target. Interestingly, the reasoning strategy employed by Coconut is not greedy search: while “lempus” initially has the highest value (0.33) at the first reasoning step (Figure 5, left), the model subsequently assigns the highest value (0.87) to “rorpus,” a child of “grimpus,” rather than following “lempus” (Figure 5, right). This characteristic resembles a breadth-first search (BFS) approach, contrasting sharply with the greedy decoding typical of traditional CoT methods. The inherent capability of continuous representations to encode multiple candidate paths enables the model to avoid making immediate deterministic decisions. Importantly, this tree search pattern is not limited to the illustrated example, but constitutes a fundamental mechanism underlying the consistent improvement observed with larger values of $k$ in Coconut.
Figure 6: Analysis of parallelism in the first two steps of the latent tree search. The three curves in each panel depict the cumulative value of the top-1, top-2, and top-3 candidate nodes.
4.4 Why is Latent Space Better for Planning?
Building upon the tree search perspective, we further examine why latent reasoning benefits planning tasks—specifically, why maintaining multiple candidate paths and postponing deterministic decisions enhances reasoning performance. Our hypothesis is that nodes explored in the early reasoning stages are inherently more challenging to evaluate accurately because they are farther from the final target nodes. In contrast, nodes positioned closer to potential targets, having fewer subsequent exploration possibilities, can be assessed accurately with higher confidence.
Figure 7: The correlation between the predicted value of correct/incorrect nodes and their heights.
To systematically test this, we define the height of a node as its shortest distance to any leaf node and analyze the relationship between node height and the model’s estimated value. Ideally, a correct node—one that can lead to the target node—should receive a high estimated value, whereas an incorrect node—one that cannot lead to the target node—should receive a low value. Empirical results across the test set (Figure 7) support our hypothesis: nodes with lower heights consistently receive more accurate and definitive probability evaluations. Conversely, nodes with greater heights exhibit more ambiguous evaluations, reflecting increased uncertainty.
These findings underscore the advantage of latent space reasoning. By delaying deterministic decisions and allowing exploration to proceed toward terminal states, latent reasoning significantly enhances the model’s ability to differentiate correct paths from incorrect ones, thereby improving performance on complex, planning-intensive tasks compared to traditional greedy methods.
5 Empirical Results with Coconut
After analyzing the promising parallel search pattern of Coconut, we validate the feasibility of LLM reasoning in a continuous latent space through more comprehensive experiments, highlighting its better reasoning efficiency over language space, as well as its potential to enhance the model’s expressivity with test-time scaling.